I put $10000.00 in the bank. In 12 years the principle and interest is $16000.00. What is the simple interest rate.
I am sure the answer is 5.5% but I do not know how to get it through the correct formula.
Thank you for your help

Answer by mananth(16949)

P=10000
Interest = 6000
n= number of years = 12
I= p*r*n
6000= 10000*r*12
6000/(10000*12)=r
6000/120000=r
6/120
=0.05
0.05 * 100 = 5%